# Designing for Novice Debuggers: A Pilot Study on an AI-Assisted Debugging Tool

### v1

@@ Abstract.
Debugging is a fundamental skill that novice programmers must develop. Numerous tools have been created to assist novice programmers in this process. Recently, large language models (LLMs) have been integrated with automated program repair techniques to generate fixes for students’ buggy code. However, many of these tools foster an over-reliance on AI and do not actively engage students in the debugging process. In this work, we aim to design an intuitive debugging assistant, CodeHinter, that combines traditional debugging tools with LLM-based techniques to help novice debuggers fix semantic errors while promoting active engagement in the debugging process. We present findings from our second design iteration, which we tested with a group of undergraduate students. Our results indicate that the students found the tool highly effective in resolving semantic errors and significantly easier to use than the first version. Consistent with our previous study, error localization was the most valuable feature. Finally, we conclude that any AI-assisted debugging tool should be personalized based on user profiles to optimize their interactions with students.
@@
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: 25th Koli Calling International Conference on Computing Education Research; November 11–16, 2025; Koli, Finlandccs: Applied computing Educationccs: Software and its engineering Software testing and debuggingccs: Social and professional topics Software engineering education
@@ 1. Introduction
Recent advances in AI-assisted programming have introduced large-language-model (LLM) tools as potential debugging assistants, capable of automatically identifying and fixing errors. LLMs have already shown strong accuracy in correcting syntax errors, opening up new possibilities for automated debugging support (Zhang et al., 2023b; Joshi et al., 2023). Moreover, LLM-based tools such as ChatGPT and GitHub Copilot can provide explanations of syntax errors along with code examples that help learners understand and correct their mistakes (OpenAI, 2022; GitHub, 2023).
@@ 2. Related Work
Although CodeAid does not directly provide solutions, it offers step-by-step guidance on necessary changes, potentially reducing the cognitive effort required for debugging. Carver and Risinger developed a traditional framework for training students in debugging (Carver and Risinger, 1987). The process begins with testing a program, followed by answering a sequence of diagnostic questions, including “what is the problem?” and “what type of bug could cause the problem?”. Katz and Anderson observed that programmers often employ backward reasoning when debugging their own code, typically starting by examining the output (Katz and Anderson, 1987). Additionally, research indicates that the skills required for fixing errors are not necessarily related to the methods used to identify and locate them. Studies show that for most students, the primary challenge in debugging lies not in repairing errors, but in the earlier stages of troubleshooting, including understanding the system, testing, and locating errors within the system (McCauley et al., 2008). In fact, a multi-institutional study of novice debuggers by Fitzgerald et al. found that once students successfully identify and locate bugs, they are generally able to fix them (Fitzgerald et al., 2008).
@@ 3. Design of CodeHinter
{subfigure}
0.45 Refer to caption {subfigure}0.45 Refer to caption
Figure 1. (Left) ‘End-to-End Test’ feature, the main feature of CodeHinter. The screenshot shows the state when users encounter failed test cases (Figure 2, marked by *), highlighting the expected value and actual output on the text editor while providing a brief explanation in the chatbot. (Right) CodeHinter ‘Expand Menu’, where users can access other features.
@@
Figure 2. Flow of users interaction with our tool. The process begins with users press the single ‘End-to-End Test’ button, which checks for syntax and semantic errors. Users are then guided to complete the missing debugging steps. If test cases fail, the tool provides options to help users resolve the bugs. The asterisk indicates the state of the screenshot in Figure 1.
@@
The helper tools in the end-to-end tests can also be manually selected by users using the main menu item ‘Helpers to Solve Bugs’. Four tools are currently provided: identifying lines with errors; providing hints and quizzes; inserting print statements; and opening Python Tutor (Guo, 2021). These features are designed to assist users in debugging their code without directly providing solutions from an LLM. They can be triggered automatically when test cases fail, or manually by expanding the menu and selecting the desired option (Figure 1). We expand on the helper tools below.
@@
Open Python Tutor. Given that novices rarely use debugger mode, to bridge this gap, the final helper tool is a button that redirects users to Python Tutor, allowing them to visualize code execution in an interactive environment. This approach provides a more intuitive and accessible debugging experience without requiring user to step into the complexity of debugger mode. We see this as a stepping stone for users, helping them gradually transition to advanced debugging techniques commonly used by more experienced programmers.
@@ 4. Methodology
After completing the tasks, participants were required to submit a post-task survey and a standardized Brooke’s system usability questionnaire (Brooke, 1996). The post-task survey consisted of common questions applicable to both CodeHinter and SID, followed by tool-specific questions. The common section assessed general usability, effectiveness, and user preferences, while the tool-specific sections focused on unique features, user confidence in debugging, and areas for improvement. Additionally, participants provided qualitative feedback on what they liked, disliked, and suggested features for future enhancements. As a token of appreciation, participants received USD 20 upon completion of the study.
@@ 5.1. Profile of Participants
We recruited eight first-year and two second-year undergraduate students from the Singapore University of Technology and Design (SUTD). The first-year students had completed an introductory programming course but had not yet chosen their major, while the second-year students may have taken additional programming courses. We found that seven out of ten participants had experience writing over 500 lines of code for unique projects, suggesting they were not complete beginners. However, their confidence levels in debugging varied. While 50% felt confident in debugging simpler problems, only 40% were comfortable handling more complex issues. Regarding debugging preferences, 60% preferred guidance that included explanations of key actions needed to identify the bug. Notably, only two students regularly used debugger tools as part of their programming habits, while five students had never heard of any debugging tools.
@@ 5.2. Post-Task Survey Results
Figure 3. Users perspective on the usefulness of CodeHinter’s features in fixing semantic errors.:
@@
As expected, ‘End-to-End Test’ was used by all participants and ranked as the most frequently used feature. Among the four available helper tools, ‘Locate Lines With Errors’ was the most popular, with most participants using it multiple times—except for one user. The ‘Provide Hint and Quiz’ and ‘Insert Print Statements’ features were each used by half of the participants. However, participants tended to use ‘Provide Hint and Quiz’ multiple times, whereas ‘Insert Print Statements’ was typically used only once per user. Additionally, only one user utilized the ‘Open Python Tutor in Web Browser’ feature, while two users interacted with the general chatbot. Notably, six out of ten participants clicked the ‘Provide Code Solution’ button, with some clicking it multiple times, possibly to check and compare their answers.
@@ 5.3.2. Locate Line With Error
“I dont have to search and debug/print a million lines to find the root cause of the error.”
@@ 5.3.3. Hint and Quiz
This feature was also well-received by users. The majority of the participants appreciated its interactive approach, as it encourages thinking when fixing the bugs. However, one comment suggested that this feature may not be their preferred option if given a limited time in solving the problems.
“The hints provided makes me think about where the error is in my code by myself.”
@@ 5.3.4. Insert Print Statement
The results for ‘Insert Print Statement’ are similar to the ‘Provide Hint and Quiz’ feature, with half of the participants attempted to use it. However, each user only used this feature once, unlike ‘Provide Hint and Quiz’, which participants interacted with multiple times. This difference may be because locations and variables of the print statements tend to be the same depending on the structure of the code. On the other hand, the ‘Provide Hint and Quiz’ feature generates different options as the code changes. This could happen if the users apply a wrong fix. Users also mentioned that Insert Print Statement helps them debug faster by allowing them to quickly spot errors, making it especially beneficial for beginners.
@@ 6. Discussion & Future Work
Regarding the ‘Insert Print Statement’ feature, we were initially surprised that each user used it only once for the problem they work with CodeHinter. However, their comments were largely positive, indicating that they found the feature useful. Therefore, we hypothesize that, unlike ‘Provide Hints and Quiz’, this feature serves a one-time function in helping users identify bugs. Once the print statements are inserted, users do not need to reinsert them; instead, they simply run the code to observe the output. Additionally, this feature was introduced as part of our design exploration, serving as an alternative to automatically inserting breakpoints, as implemented in the first iteration of our tool, SID. Our initial hypothesis was that novice programmers find using print statements easier than working with an IDE’s debug mode. However, we discovered that some intermediate programmers prefer using breakpoints and the debug mode instead. This suggests that future iterations of the tool should offer both options, allowing users to choose the method that best suits their debugging preferences.
@@ 7. Conclusion
In conclusion, this study demonstrates the potential of AI-assisted tools like CodeHinter to enhance the debugging experience for novice programmers by fostering active engagement rather than passive reliance on automated solutions. Through a user-centric design that integrates spectrum-based fault localization, interactive hints and quizzes, and print statements suggestions, all within a familiar IDE environment, CodeHinter effectively supports the debugging process. The tool’s usability and effectiveness were validated in a pilot study, with participants favouring features that guided them through the problem-solving process while reinforcing critical thinking. Our findings also underscore the importance of tailoring AI support to individual user profiles and problem contexts. Future debugging tools should focus on providing personalized assistance to optimize learning outcomes and foster independent debugging skills.
8. Acknowledgments

### v2

@@ Abstract.
Debugging is a fundamental skill that novice programmers must develop. Numerous tools have been created to assist novice programmers in this process. Recently, large language models (LLMs) have been integrated with automated program repair techniques to generate fixes for students’ buggy code. However, many of these tools foster an over-reliance on AI and do not actively engage students in the debugging process. In this work, we aim to design an intuitive debugging assistant, CodeHinter, that combines traditional debugging tools with LLM-based techniques to help novice debuggers fix semantic errors while promoting active engagement in the debugging process. We present findings from our second design iteration, which we tested with a group of undergraduate students. Our results indicate that the students found the tool highly effective in resolving semantic errors and significantly easier to use than the first version. Consistent with our previous study, error localization was the most valuable feature. Finally, we conclude that any AI-assisted debugging approach should be personalized based on user profiles to optimize their interactions with the tool.
@@
journalyear: 2025copyright: ccconference: 25th Koli Calling International Conference on Computing Education Research; November 11–16, 2025; Koli, Finlandbooktitle: 25th Koli Calling International Conference on Computing Education Research (Koli Calling ’25), November 11–16, 2025, Koli, Finlanddoi: 10.1145/3769994.3769997isbn: 979-8-4007-1599-0/25/11ccs: Applied computing Educationccs: Software and its engineering Software testing and debuggingccs: Social and professional topics Software engineering education
@@ 1. Introduction
Recent advances in AI-assisted programming have introduced large language model (LLM) tools as potential debugging assistants, capable of automatically identifying and fixing errors. LLMs have already shown strong accuracy in correcting syntax errors, opening up new possibilities for automated debugging support (Zhang et al., 2023b; Joshi et al., 2023). Moreover, LLM-based tools such as ChatGPT and GitHub Copilot can provide explanations of syntax errors along with code examples that help learners understand and correct their mistakes (OpenAI, 2022; GitHub, 2023).
@@ 2. Related Work
Although CodeAid does not directly provide solutions, it offers step-by-step guidance on necessary changes, potentially reducing the cognitive effort required for debugging. Carver and Risinger developed a traditional framework for training students in debugging (Carver and Risinger, 1987). The process begins with testing a program, followed by answering a sequence of diagnostic questions, including “what is the problem?” and “what type of bug could cause the problem?”. Katz and Anderson observed that programmers often employ backward reasoning when debugging their own code, typically starting by examining the output (Katz and Anderson, 1987). Additionally, research indicates that the skills required for fixing errors are not necessarily related to the methods used to identify and locate them. Studies show that for most students, the main difficulty in debugging lies not in repairing errors, but in earlier troubleshooting stages, such as understanding the system, testing, and locating errors within it (McCauley et al., 2008). In fact, a multi-institutional study of novice debuggers by Fitzgerald et al. found that once students successfully identify and locate bugs, they are generally able to fix them (Fitzgerald et al., 2008).
@@ 3. Design of CodeHinter
Figure 1. (Left) ‘End-to-End Test’ feature, the main feature of CodeHinter. The screenshot shows the state when users encounter failed test cases (corresponding to * in Figure 2), highlighting the expected value and actual output in the text editor while providing a brief explanation in the chatbot. (Right) CodeHinter ‘Expand Menu’, where users can access other features.
@@
Figure 2. Flow of user interactions with our tool. The process begins with users pressing the ‘End-to-End Test’ button, which checks for syntax and semantic errors. Users are then guided to complete the missing debugging steps. If test cases fail, the tool provides options to help users resolve the bugs. The asterisk indicates the state of the screenshot in Figure 1.
@@
The helper tools in the end-to-end tests can also be manually selected by using the main menu item ‘Helpers to Solve Bugs’. Four tools are currently provided: identifying lines with errors; providing hints and quizzes; inserting print statements; and opening Python Tutor (Guo, 2021). These features are designed to assist users in debugging their code without directly providing solutions from an LLM. They can be triggered automatically when test cases fail, or manually by expanding the menu and selecting the desired option (Figure 1). We expand on the helper tools below.
@@
Open Python Tutor. Given that novices rarely use debugger mode, to bridge this gap, the final helper tool is a button that redirects users to Python Tutor, allowing them to visualize code execution in an interactive environment. This approach provides a more intuitive and accessible debugging experience without requiring users to step into the complexity of debugger mode. We see this as a stepping stone for users, helping them gradually transition to advanced debugging techniques commonly used by more experienced programmers.
@@ 4. Methodology
After completing the tasks, participants were required to submit a post-task survey and a standardized Brooke’s system usability questionnaire (Brooke, 1996). The post-task survey consisted of common questions applicable to both CodeHinter and SID, followed by tool-specific questions. The common section assessed general usability, effectiveness, and user preferences, while the tool-specific sections focused on unique features, user confidence in debugging, and areas for improvement. Additionally, participants provided qualitative feedback on what they liked, disliked, and suggested features for future enhancements. As a token of appreciation, participants received the equivalent of USD 20 upon completion of the study.
@@ 5.1. Profile of Participants
We recruited eight first-year and two second-year undergraduate students from the Singapore University of Technology and Design (SUTD). The first-year students had completed an introductory programming course but had not yet chosen their major, while the second-year students may have taken additional programming courses. We found that seven out of ten participants had experience writing over 500 lines of code for unique projects, suggesting they were not complete beginners. However, their confidence levels in debugging varied. While 50% felt confident in debugging simpler problems, only 40% were comfortable handling more complex issues. Regarding debugging preferences, 60% preferred guidance that included explanations of key actions needed to identify the bug. Notably, only two students regularly used debugging tools as part of their programming habits, while five students had never heard of any debugging tools.
@@ 5.2. Post-Task Survey Results
Figure 3. User perspectives on the usefulness of CodeHinter’s features for fixing semantic errors.:
@@
As expected, ‘End-to-End Test’ was used by all participants and ranked as the most frequently used feature. Among the four available helper tools, ‘Locate Lines With Errors’ was the most popular, with most participants using it multiple times, except for one user. The ‘Provide Hint and Quiz’ and ‘Insert Print Statements’ features were each used by half of the participants. However, participants tended to use ‘Provide Hint and Quiz’ multiple times, whereas ‘Insert Print Statements’ was typically used only once per user. Additionally, only one user utilized the ‘Open Python Tutor in Web Browser’ feature, while two users interacted with the general chatbot. Notably, six out of ten participants clicked the ‘Provide Code Solution’ button, with some clicking it multiple times, possibly to check and compare their answers.
@@ 5.3.2. Locate Line With Error
“I don’t have to search and debug/print a million lines to find the root cause of the error.”
@@ 5.3.3. Hint and Quiz
This feature was also well-received by users. The majority of the participants appreciated its interactive approach, as it encourages thinking when fixing the bugs. However, one comment suggested that this feature may not be their preferred option if given a limited time to solve the problems in.
“The hints provided make me think about where the error is in my code by myself.”
@@ 5.3.4. Insert Print Statement
The results for ‘Insert Print Statement’ are similar to the ‘Provide Hint and Quiz’ feature, with half of the participants attempting to use it. However, each user only used this feature once, unlike ‘Provide Hint and Quiz’, which participants interacted with multiple times. This difference may be because locations and variables of the print statements tend to be the same depending on the structure of the code. On the other hand, the ‘Provide Hint and Quiz’ feature generates different options as the code changes. This could happen if the users apply a wrong fix. Users also mentioned that ‘Insert Print Statement’ helps them debug faster by allowing them to quickly spot errors, making it especially beneficial for beginners.
@@ 6. Discussion & Future Work
Regarding the ‘Insert Print Statement’ feature, we were initially surprised that each user used it only once for the problem they worked on with CodeHinter. However, their comments were largely positive, indicating that they found the feature useful. Therefore, we hypothesize that, unlike ‘Provide Hints and Quiz’, this feature serves a one-time function in helping users identify bugs. Once the print statements are inserted, users do not need to reinsert them; instead, they simply run the code to observe the output. Additionally, this feature was introduced as part of our design exploration, serving as an alternative to automatically inserting breakpoints, as implemented in the first iteration of our tool, SID. Our initial hypothesis was that novice programmers find using print statements easier than working with an IDE’s debug mode. However, we discovered that some intermediate programmers prefer using breakpoints and the debug mode instead. This suggests that future iterations of the tool should offer both options, allowing users to choose the method that best suits their debugging preferences.
@@ 7. Conclusion
In conclusion, this study demonstrates the potential of AI-assisted tools like CodeHinter to enhance the debugging experience for novice programmers by fostering active engagement rather than passive reliance on generating complete solutions. Through a user-centric design that integrates spectrum-based fault localization, interactive hints and quizzes, and print statement suggestions, all within a familiar IDE environment, CodeHinter effectively supports the debugging process. The tool’s usability and effectiveness were validated in a pilot study, with participants favoring features that guided them through the problem-solving process while reinforcing critical thinking. Our findings also underscore the importance of tailoring AI support to individual user profiles and problem contexts. Future debugging tools should focus on providing personalized assistance to optimize learning outcomes and foster independent debugging skills.
Acknowledgements.